OpenStudy (anonymous):

Can someone tell me what I'm doing wrong here? "Write the expression as a sum and/or difference of logarithms. Express powers as factors." log3((x^9)/(x-6)) , x > 6 And I answered "9log3x - 3log(x-6)" and it was incorrect. How do I fix it? (Rewriting to show clearer problem in comments.)

OpenStudy (anonymous):

\[\log _{3}(x ^{9}/(x-6)), x > 6\] <-- Problem \[9\log _{3}x - 3 \log (x-6)\] <--- My incorrect answer.

OpenStudy (shiraz14):

It should be 9log3(x) - log3(x-6). You mistreated the base as a power in your answer (only powers can be brought to the front of a log function, the base defines the log function and cannot be manipulated unless using a base-change).

OpenStudy (anonymous):

So to re-iterate, I just need to move my 3 to the other side of the 2nd log?

OpenStudy (shiraz14):

The '3' is the base of the log in this case, i.e. to say, you are using log3 of a variable.

OpenStudy (anonymous):

I see now, thank you. :)

OpenStudy (shiraz14):

To make it simpler to understand, consider the following: 3^x = y If we wanted to find out what x is, we will use the following identity: \[x = \log_{3}y\]

OpenStudy (shiraz14):

You're welcome